OpenStudy (anonymous):

Solve the equation by factoring: 2x^2-6x=0 x= x=

OpenStudy (cwrw238):

2x(x - 3) = 0

OpenStudy (anonymous):

First factor: 2x (x - 3) = 0 Next, set each term equal to zero (you can do this because they are multiplied). So, 2x = 0 x = 0 / 2 x = 0 AND x - 3 = 0 x = 3 SO, x = 0 AND/OR x = 3 You can also check these solutions by plugging them back into the original equation.